OpenStudy (anonymous):

In a stable there are men and horses. In all,there are 22 heads and 72 feet. How many men and horses are in the stable ?

OpenStudy (anonymous):

let no. of men =x and no. of horses=y x+y=22 2x+4y=72 or x+2y=36 solve for x and y

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

This ones kind of hard

OpenStudy (anonymous):

Im gonna leave this to the genius my mind just got blown

OpenStudy (anonymous):

x+2y=36 x+y+y=36 22+y=36 y=36-22=? x=22-y=?

OpenStudy (anonymous):

x+2y=36 x+y+y=36 22+y=36 y=? then x=22-y